OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (2, -2) and (-1, 10)

OpenStudy (anonymous):

y - 10 = -4(x +1)

OpenStudy (dumbcow):

agreed move the -10 over y = -4x +6

OpenStudy (amistre64):

2 -2 1 -10 ------ 3 -12; m = -12/3 = -4 y = -4x -(-4.2)-2 y = -4x +8-2 y = -4x +6